Find the output voltage of this zener diode source
Answer #1

(a) From the data sheet available, voltage of Zener diode (1N5231B) is 5.1 volt.

(b) Since voltage is 5.1 volt hence current will be

I=\frac{V}{R}=\frac{5.1}{20\times10^3}=0.255$ mA$

(c) & (d) data sheet is not available hence cant find the solution